# An algebraic theory of Łojasiewicz exponents

### v1

@@ 12.3. Uniformity in families and wall–chamber decomposition
Theorem 12.3 (Polyhedral wall–chamber structure).
@@
${\mathfrak{a}_{\bullet}}(s)=\{\mathfrak{a}(s)_{t}\}_{t\geq 0},\quad{\mathfrak{b}_{\bullet}}(s)=\{\mathfrak{b}(s)_{t}\}_{t\geq 0},$
@@
Then, there exists a finite polyhedral decomposition $S=\bigcup_{\alpha}C_{\alpha}$ such that on each chamber $C_{\alpha}$ there is an index $i(\alpha)\in\{1,\dots,N\}$ with
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}\big({\mathfrak{a}_{\bullet}}(s)\big)\ =\ \frac{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}\big({\mathfrak{a}_{\bullet}}(s)\big)}{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}\big({\mathfrak{b}_{\bullet}}(s)\big)}\quad\text{for all }s\in C_{\alpha}.$
In particular, $s\mapsto\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))$ is piecewise rational-linear and the maximizing divisor is locally constant on chambers.
@@
Fix the common model $\pi:Y\to\operatorname{Spec}R$ as in Theorem 12.2 and its exceptional prime divisors $E_{1},\dots,E_{N}$. Write $E=\sum_{i=1}^{N}E_{i}$.
Step 1: linearity of divisor coefficients. By Theorem 12.2, for each $s$ we can write
$\mathfrak{a}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\big(-D^{(s)}_{\mathfrak{a}}(t)\big),\quad\mathfrak{b}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\big(-D^{(s)}_{\mathfrak{b}}(t)\big),$
@@
$D^{(s)}_{\mathfrak{a}}(t)=\sum_{i=1}^{N}d^{(s)}_{\mathfrak{a},i}(t)\,E_{i},\quad D^{(s)}_{\mathfrak{b}}(t)=\sum_{i=1}^{N}d^{(s)}_{\mathfrak{b},i}(t)\,E_{i}$
and each coefficient $d^{(s)}_{\mathfrak{a},i}(t)$ (resp. $d^{(s)}_{\mathfrak{b},i}(t)$) is an integer-valued piecewise linear function whose slope in $t$ is a piecewise linear function of $s$. Hence, for each fixed $i$ the asymptotic slopes
$\mu_{\mathfrak{a},i}(s):=\lim_{t\to\infty}\frac{1}{t}\operatorname{\mathrm{ord}}_{E_{i}}\big(\mathfrak{a}(s)_{t}\big),\quad\mu_{\mathfrak{b},i}(s):=\lim_{t\to\infty}\frac{1}{t}\operatorname{\mathrm{ord}}_{E_{i}}\big(\mathfrak{b}(s)_{t}\big)$
exist and are piecewise linear functions of $s$.
Step 2: finite-max formula on the fixed testing set. By the finite-testing criterion, Theorem 5.8, with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$ (we can restrict to those with $\mu_{\mathfrak{b},i}(s)>0$),
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))\ =\ \max_{1\leq i\leq N}\ \frac{\mu_{\mathfrak{a},i}(s)}{\mu_{\mathfrak{b},i}(s)}.$
For each ordered pair $(i,j)$, the inequality
$\frac{\mu_{\mathfrak{a},i}(s)}{\mu_{\mathfrak{b},i}(s)}\ \geq\ \frac{\mu_{\mathfrak{a},j}(s)}{\mu_{\mathfrak{b},j}(s)}$
is equivalent to the linear inequality
$\mu_{\mathfrak{a},i}(s)\mu_{\mathfrak{b},j}(s)-\mu_{\mathfrak{a},j}(s)\mu_{\mathfrak{b},i}(s)\ \geq\ 0.$
By Theorem 12.2, for each fixed exceptional divisor $E_{i}$ the functions
@@
$\mu_{\mathfrak{a},i}(t)=\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{a}_{t})\quad\text{and}\quad\mu_{\mathfrak{b},i}(s)=\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{b}_{s})$
are piecewise affine linear functions of $t$ and $s$, respectively. More precisely, there exist real constants $c_{\mathfrak{a},i},c_{\mathfrak{b},i}$ and bounded error terms $\varepsilon_{\mathfrak{a},i}(t),\varepsilon_{\mathfrak{b},i}(s)$ such that
$\mu_{\mathfrak{a},i}(t)=c_{\mathfrak{a},i}\,t+\varepsilon_{\mathfrak{a},i}(t),\quad\mu_{\mathfrak{b},i}(s)=c_{\mathfrak{b},i}\,s+\varepsilon_{\mathfrak{b},i}(s),$
where $\varepsilon_{\mathfrak{a},i}(t)$ (resp. $\varepsilon_{\mathfrak{b},i}(s)$) is locally constant away from a finite set of jump points determined by the ceiling functions $\lceil\lambda_{j}t\rceil$ (resp. $\lceil\lambda_{j}s\rceil$). Consequently, on each region of the (s,t)(s,t)–plane where these ceiling functions are fixed, the functions $\mu_{\mathfrak{a},i}(t)$ and $\mu_{\mathfrak{b},i}(s)$ are affine linear. Hence, refining $S$ by the finite arrangement of the linearity regions of all $\mu_{\mathfrak{a},i},\mu_{\mathfrak{b},i}$ yields a finite polyhedral decomposition on which all comparisons are given by linear inequalities.
Step 3: chambers and constancy of maximizers. Intersecting the regions defined by the finitely many pairwise comparison inequalities for the index attaining the maximum produces the claimed chambers $C_{\alpha}$, and on each chamber the maximizing index $i(\alpha)$ is constant. ∎
@@ 13. Concluding remarks and open problems
Set $I=\mathfrak{a}^{p}$. We use the following one–function statement, which is a specialization of [36, Théorème 7.2] (equivalence of conditions (1),(5),(6)):
@@
Fix such $V$ and choose $V^{\prime}\subseteq V$.
@@
Since $I=\mathfrak{a}^{p}$ is generated by finitely many degree – $p$ monomials $F_{\nu}=f_{1}^{e_{1,\nu}}\cdots f_{r}^{e_{r,\nu}}$, their pullbacks generate $(s)$. Hence, there exist holomorphic $v_{\nu}$ on $V$ such that
@@
Step 2: $(LT\text{--}S)$, first direction. Assume $h\in\overline{I}$. By [36, Théorème 7.2], there exists a neighborhood $\tilde{W}$ of $\pi^{-1}(0)$ such that
@@
Step 3: $(LT\text{--}S)$, converse direction. Assume $|h(x)|\leq C\|f(x)\|^{p}$ on a neighborhood $U$ of 0. Pulling back,
$|h\circ\pi|\leq C\|f\circ\pi\|^{p}.$
On $V^{\prime}\cap\pi^{-1}(U)$, using $\|f\circ\pi\|^{p}\leq A|s|$,
$|h\circ\pi|\leq CA|s|.$
Hence, $(h\circ\pi)/s$ is holomorphic and bounded on $(V^{\prime}\cap\pi^{-1}(U))\setminus\{s=0\}$ and extends holomorphically across $\{s=0\}$. Thus, $h\circ\pi\in(s)=I\mathcal{O}_{Y}$ near $\pi^{-1}(0)$. By Lejeune-Jalabert–Teissier, Théorème 7.2, $h\in\overline{I}$.

### v2

@@ 12.3. Uniformity in families and wall–chamber decomposition
Theorem 12.3 (Wall–chamber stratification).
@@
${\mathfrak{a}_{\bullet}}(s)=\{\mathfrak{a}(s)_{t}\}_{t\geq 0}\text{ and }{\mathfrak{b}_{\bullet}}(s)=\{\mathfrak{b}(s)_{t}\}_{t\geq 0},$
@@
Then, there exists a finite stratification $S=\bigsqcup_{\alpha}C_{\alpha}$ by locally closed semialgebraic subsets such that on each stratum $C_{\alpha}$ one of the following holds:
  1. (i)
    $\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\infty$ for all $s\in C_{\alpha}$;
  2. (ii)
    there exists a nonempty subset $J(\alpha)\subseteq\{1,\dots,N\}$ such that for every $s\in C_{\alpha}$,
    $\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\max_{i\in J(\alpha)}\frac{\operatorname{\mathrm{ord}}_{E_{i}}({\mathfrak{a}_{\bullet}}(s))}{\operatorname{\mathrm{ord}}_{E_{i}}({\mathfrak{b}_{\bullet}}(s))},$
    and the maximizer set
    $\left\{i\in J(\alpha)\ \middle|\ \frac{\operatorname{\mathrm{ord}}_{E_{i}}({\mathfrak{a}_{\bullet}}(s))}{\operatorname{\mathrm{ord}}_{E_{i}}({\mathfrak{b}_{\bullet}}(s))}=\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))\right\}$
    is independent of $s\in C_{\alpha}$.
In particular, on every stratum $C_{\alpha}$ for which the maximizer set is a singleton $\{i(\alpha)\}$, one has
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\frac{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}({\mathfrak{a}_{\bullet}}(s))}{\operatorname{\mathrm{ord}}_{E_{i(\alpha)}}({\mathfrak{b}_{\bullet}}(s))}\quad\text{for all }s\in C_{\alpha}.$
Hence, on such a stratum the function
$s\longmapsto\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))$
is given by a single ratio of affine linear functions.
@@
Fix the common model $\pi:Y\to\operatorname{Spec}R$ as in Theorem 12.2 and its exceptional prime divisors $E_{1},\dots,E_{N}$.
Step 1: affine linearity of the asymptotic divisor coefficients. By Theorem 12.2, for each $s\in S$ and each $t\geq 0$ we have
$\mathfrak{a}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\bigl(-D^{(s)}_{\mathfrak{a}}(t)\bigr)\text{ and }\mathfrak{b}(s)_{t}=\pi_{*}\mathcal{O}_{Y}\bigl(-D^{(s)}_{\mathfrak{b}}(t)\bigr),$
@@
$D^{(s)}_{\mathfrak{a}}(t)=\sum_{j}\lceil\lambda_{\mathfrak{a},j}(s)t\rceil\,F_{\mathfrak{a},j}\text{ and }D^{(s)}_{\mathfrak{b}}(t)=\sum_{j}\lceil\lambda_{\mathfrak{b},j}(s)t\rceil\,F_{\mathfrak{b},j},$
with $F_{\mathfrak{a},j}=\sum_{i=1}^{N}a_{\mathfrak{a},ij}E_{i}\text{ and }F_{\mathfrak{b},j}=\sum_{i=1}^{N}a_{\mathfrak{b},ij}E_{i}.$ Hence,
@@
$\operatorname{\mathrm{ord}}_{E_{i}}\bigl(\mathfrak{a}(s)_{t}\bigr)=\sum_{j}a_{\mathfrak{a},ij}\,\lceil\lambda_{\mathfrak{a},j}(s)t\rceil\text{ and }\operatorname{\mathrm{ord}}_{E_{i}}\bigl(\mathfrak{b}(s)_{t}\bigr)=\sum_{j}a_{\mathfrak{b},ij}\,\lceil\lambda_{\mathfrak{b},j}(s)t\rceil.$
Dividing by $t$ and letting $t\to\infty$, we obtain
$\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)=\sum_{j}a_{\mathfrak{a},ij}\lambda_{\mathfrak{a},j}(s)\text{ and }\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)=\sum_{j}a_{\mathfrak{b},ij}\lambda_{\mathfrak{b},j}(s).$
Since the coefficient functions $\lambda_{\mathfrak{a},j}(s)$ and $\lambda_{\mathfrak{b},j}(s)$ are piecewise linear, it follows that each function
[…]
is piecewise affine linear, and so continuous on each member of a finite decomposition of $S$.
Step 2: separate the infinite-valued region. Refining the above decomposition if necessary, we may assume that for each fixed $i$ the sign of $\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)$ is constant on each piece. Since these quantities are nonnegative, on each piece $P$ there is a well-defined subset
$J_{P}:=\bigl\{\,i\in\{1,\dots,N\}\mid\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)>0\text{ for all }s\in P\,\bigr\}.$
If on such a piece $P$ there exists $i\notin J_{P}$ such that $\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)>0\text{ for all }s\in P,$ then for every $s\in P$, one has
$\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)=0\text{ and }\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)>0.$
By the finite-testing criterion in Theorem 12.2, no finite slope can satisfy the containment inequalities along $E_{i}$. Hence,
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\infty\text{ for all }s\in P.$
Step 3: reduction to Theorem 9.9 on the finite-valued region. Now restrict to a piece $P$ on which no such index occurs. Then, for every $i\notin J_{P}$ one has
$\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)=0\quad\Longrightarrow\quad\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)=0\quad\text{ for all }s\in P.$
Thus, only the divisors $E_{i}$ with $i\in J_{P}$ contribute to the finite-testing formula of Theorem 12.2.
For each $s\in P$, define $\alpha_{i}(s):=\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{a}_{\bullet}}(s)\bigr)\text{ and }\beta_{i}(s):=\operatorname{\mathrm{ord}}_{E_{i}}\bigl({\mathfrak{b}_{\bullet}}(s)\bigr)\quad(i\in J_{P}).$ Then,
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))=\max_{i\in J_{P}}\frac{\alpha_{i}(s)}{\beta_{i}(s)},$
and, by construction, each $\alpha_{i}$ and $\beta_{i}$ is continuous on $P$, while $\beta_{i}(s)\in(0,\infty)\text{ for all }i\in J_{P},\ s\in P.$ Therefore, the hypotheses of Theorem 9.9 are satisfied on $P$ with finite candidate set $V_{P}:=\{\operatorname{\mathrm{ord}}_{E_{i}}\mid i\in J_{P}\}.$ Applying Theorem 9.9, we obtain a finite partition of $P$ into locally closed strata on each of which the maximizer set is constant. In particular, on each such stratum the function $s\longmapsto\mathcal{L}_{{\mathfrak{b}_{\bullet}}(s)}({\mathfrak{a}_{\bullet}}(s))$ is given by a single valuative ratio whenever the maximizer set is a singleton.
Taking the common refinement over all pieces $P$ gives the required finite stratification of $S$. ∎
@@ 13. Concluding remarks and open problems
Set $I=\mathfrak{a}^{p}$. We use the following one–function statement, which is a specialization of [40, Théorème 7.2] (equivalence of conditions (1),(5),(6)):
@@
Fix such $V$, and choose $V^{\prime}\Subset V$ a relatively compact subset of $V$.
@@
Since $I=\mathfrak{a}^{p}$ is generated by finitely many degree – $p$ monomials $F_{\nu}=f_{1}^{e_{1,\nu}}\cdots f_{s}^{e_{s,\nu}}$, their pullbacks generate $(s)$. Hence, there exist holomorphic $v_{\nu}$ on $V$ such that
@@
Step 2: $(LT\text{--}S)$, first direction. Assume $h\in\overline{I}$. Then, there exists a neighborhood $\tilde{W}$ of $\pi^{-1}(0)$ such that
@@
Step 3: $(LT\text{--}S)$, converse direction. Assume $|h(x)|\leq C\|f(x)\|^{p}$ on a neighborhood $U$ of 0. Pulling back, we get
$|h\circ\pi|\leq C\|f\circ\pi\|^{p}$
on $\pi^{-1}(U)$.
Choose finitely many relatively compact open sets $V_{1},\dots,V_{r}\Subset Y$ covering $\pi^{-1}(0)$ such that for each $\alpha$ the invertible ideal $I\mathcal{O}_{Y}|_{V_{\alpha}}$ is principal, say $I\mathcal{O}_{Y}(V_{\alpha})=(s_{\alpha}),$ with $s_{\alpha}\in\mathcal{O}_{Y}(V_{\alpha})$. Shrinking if necessary, we may assume that each $V_{\alpha}$ is contained in one of the neighborhoods $V$ used in Step 1. Hence, for each $\alpha$, there exists a constant $A_{\alpha}>0$ such that on $V_{\alpha}$ one has $\|f\circ\pi\|^{p}\leq A_{\alpha}|s_{\alpha}|.$ Therefore, on $(V_{\alpha}\cap\pi^{-1}(U))\setminus\{s_{\alpha}=0\}$,
$|h\circ\pi|\leq C\|f\circ\pi\|^{p}\leq CA_{\alpha}|s_{\alpha}|,$
so
$\left|\frac{h\circ\pi}{s_{\alpha}}\right|\leq CA_{\alpha}.$
Since $Y$ is normal and $s_{\alpha}$ is a local generator of the invertible ideal $I\mathcal{O}_{Y}|_{V_{\alpha}}$, the germ of $s_{\alpha}$ is a non-zero-divisor at every point of $V_{\alpha}$. Thus $(h\circ\pi)/s_{\alpha}$ defines a meromorphic function on $V_{\alpha}\cap\pi^{-1}(U)$, and the boundedness above implies that it is holomorphic there. Hence
$h\circ\pi\in s_{\alpha}\mathcal{O}_{Y}(V_{\alpha}\cap\pi^{-1}(U))=I\mathcal{O}_{Y}(V_{\alpha}\cap\pi^{-1}(U))$
for every $\alpha$.
Since the sets $V_{\alpha}$ cover $\pi^{-1}(0)$, it follows that $h\circ\pi\in I\mathcal{O}_{Y}$ on a neighborhood of $\pi^{-1}(0)$. Let $E$ be any exceptional divisor of the normalized blow-up. Then
$\operatorname{\mathrm{ord}}_{E}(h\circ\pi)\geq\operatorname{\mathrm{ord}}_{E}(I\mathcal{O}_{Y}).$
These divisors correspond to the Rees valuations of $I$, so the above inequalities imply $v(h)\geq v(I)$ for every such valuation $v$. Hence, $h\in\overline{I}$ by Lemma 5.2.